# Sharp non-uniqueness for the Navier–Stokes equations in scaling critical spaces

### v1

@@ 1.3. Outline of the proof
Let us explain the strategy of the proof of Theorem 1.2. The strategy is that we first build a dominant part of the solution and then solve the remaining perturbative problem by a fixed point argument in a function space where well-posedness is available. More precisely, we construct an almost frequency-localized bulding blocks $\{V_{j}\}_{j\geqslant 0}$ of stationary flows, each essentially supported on the $j$th frequency shell, and design the principal profile
$V=\sum_{j=0}^{\infty}V_{j}\neq 0\qquad{\rm in}\quad\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})$ (1.10)
@@
$F:=-\Delta V+\mathbb{P}\operatorname{div}(V\otimes V),$ (1.11)
@@
${\left\|W\right\|}_{\dot{B}_{r,1}^{n/r-1}}\lesssim{\left\|F\right\|}_{\dot{B}_{r,1}^{n/r-3}}\ll{\left\|V\right\|}_{\dot{B}_{p,q}^{n/p-1}}.$ (1.13)
Hence, the desired steady state is obtained in the form $U=V+W$. Note that ${\left\|U\right\|}_{\dot{B}_{p,q}^{n/p-1}}\sim{\left\|V\right\|}_{\dot{B}_{p,q}^{n/p-1}}\neq 0$ holds. Furthermore, to construct a non-stationary solution $u$ whose asymptotic profile is $U$, we decompose $u=U+w$. Then, we may follow the standard stability argument to construct a small dissipative perturbed flow $w\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some $2<\theta<\infty$. Then, the auxilliary dissipative space enables us to show $\|w(t)\|_{\dot{B}_{p,q}^{n/p-1}}\to 0$ as $t\to\infty$.
We now explain how to construct the leading profile $V$ satisfying (1.12). A natural first attempt, motivated by existing works such as [Coi-Pal-25], would be to choose each $V_{j}$ as a superposition of plane waves concentrated on a single highly lacunary frequency scale $N_{j}=A^{\gamma b^{j}}$ ($A\gg 1$), for instance, $V_{j}(x)=\sum_{k\in\mathcal{K}}f_{k,j}(x)\cos(N_{j}a_{k}\cdot x)k$, where $\mathcal{K}\subset\mathbb{R}^{n}$ is a finite set, $f_{k,j}$ are smooth amplitudes, and $a_{k}$ are constant vectors with $a_{k}\cdot k=0$. However, such a single-scale construction is too rigid to capture the delicate dependence on the Besov summability index $q$, which is essential for observing the subtle non-uniqueness phenomena addressed in (N2). Instead, we distribute the $j$th building block over many internal scales. To this end, for $\lambda\gg 1$ (which corresponds to $\varepsilon^{-1/\alpha}$ with some $\alpha$) and a fixed large number $\mu\in\mathbb{N}$, we introduce a tetration-like sequence [^4] This is defined by $\lambda_{0}=\lambda$ and $\lambda_{j}=\lambda^{2^{\lambda_{j-1}^{4\mu}}}$; see (4.9). $\{\lambda_{j}\}_{j\geqslant 0}$ and define the $j$th level frequency index set by
@@
and take $V_{j}$ in the schematic form
@@
Here, the choice of $1/\sqrt{h_{j}\ell}$ is inspired by [Tsu-19-ARMA]. With this multi-scale design, the oscillation calculus in Section 3 below yields
@@
The cross-shell interaction term $F_{2}$ is favorable, as it consists of products of flows concentrated on distinct frequency shells. By contrast, $F_{1}$ exhibits a more delicate structure. Indeed, within the nonlinear interaction $\operatorname{div}(V_{j}\otimes V_{j})$, resonances among modes inside the frequency shell $\Lambda_{j}$ generate an inverse-cascade transfer toward low frequencies, which exhibit the worst regularity. By selecting the directional set $\mathcal{K}$ and the amplitudes $f_{k,j}$ appropriately through Nash’s geometric lemma, one can arrange for this low-frequency contribution to cancel $\Delta V_{j-1}$. Consequently, the forcing term $F$ attains improved regularity and smaller magnitude relative to $V$.
@@ 2. Preliminaries
Throughout this paper, we denote by $C$ and $c$ the constants, which may differ in each line. In particular, C=C(,,)C=C(*,...,*) denotes the constant which depends only on the quantities appearing in parentheses. Let $\mathbb{N}$ be the set of all positive integers and put $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. For $\mathbb{M}\subset\mathbb{Z}$, we define $2^{\mathbb{M}}:=\{2^{m}\ ;\ m\in\mathbb{M}\}$. The identity matrix is denoted by $\rm Id$}. For a metric space $(X,d)$, $x_{0}\in X$, and a positive number $r>0$, set $B_{X}(x_{0},r):=\{x\in X\ ;\ d(x,x_{0})\leqslant r\}$. Let $\mathscr{S}(\mathbb{R}^{n})$ be the set of all Schwartz functions on $\mathbb{R}^{n}$, and we denote by $\mathscr{S}^{\prime}(\mathbb{R}^{n})$ the set of all tempered distributions on $\mathbb{R}^{n}$.
@@
$\Delta_{j}f:=\mathscr{F}^{-1}\left[\widehat{\varphi_{j}}(\xi)\widehat{f}(\xi)\right],\qquad f\in\mathscr{S}^{\prime}(\mathbb{R}^{n}),\quad j\in\mathbb{Z}.$ (2.3)
For $1\leqslant p,q\leqslant\infty$ and $s\in\mathbb{R}$, the Besov space $\dot{B}_{p,q}^{s}(\mathbb{R}^{n})$ is defined as
@@
there exists a positive constant $C=C(n,p,q,r)$ such that
@@
The Chemin–Lerner spaces are well-known to be particularly convenient when exploiting the maximal regularity of the heat kernel, and we take advantage of this framework through the following lemma, which will be used in our stability analysis for non-stationary flows around the steady state.
@@
Let $n\geqslant 2$. Let $1\leqslant p,q,\theta\leqslant\infty$ satisfy
$2<\theta<\infty,\qquad\min\left\{n,\frac{2n}{p}\right\}-2+\frac{2}{\theta}>0.$ (2.12)
Then, there exists a positive constant $C=C(n,p,q,\theta)$ such that
@@
for all $I=(t_{0},t_{1})\subset\mathbb{R}$, $v\in\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))$, and $w\in\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}}(\mathbb{R}^{n}))$.
@@ 3. Highly oscillating functions in Besov spaces
In the proof of our main result, we frequently use the sharp Besov estimates for finite superpositions of functions obtained by modulating a smooth amplitude with a plane wave of the form (3.4) below. If each Fourier transform $\widehat{f_{k}}$ of the amplitudes were compactly supported in a low frequency region, then the Fourier support of each $g_{\lambda,\ell,k}$ is concentrated around $\xi\sim\lambda^{\ell}a_{k}$, and we have
${\left\|g_{\lambda,\ell,k}\right\|}_{\dot{B}_{p,q}^{s}}\sim|b_{\ell}|(\lambda^{\ell}|a_{k}|)^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.1)
@@
$\|F_{\lambda,L}\|_{\dot{B}^{s}_{p,q}}\sim\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}.$ (3.2)
@@
Let $K,L\in\mathbb{N}$. Let $a_{1},\dots,a_{K}\subset\mathbb{R}^{n}\setminus\{0\}$ satisfy $|a_{k}|\in 2^{\mathbb{N}}$ $(k=1,\dots,K)$ and assume that $a_{k}$, $a_{k^{\prime}}$ do not point in the same direction unless $k=k^{\prime}$. For $\lambda\in 2^{\mathbb{N}}$, $b_{1},\dots,b_{L}\in\mathbb{R}$, and $f_{1},\dots,f_{K}\in\mathscr{S}(\mathbb{R}^{n})$, set
$F_{\lambda,L}(x):=\sum_{k=1}^{K}\sum_{\ell=1}^{L}g_{\lambda,\ell,k}(x),\qquad g_{\lambda,\ell,k}(x):=b_{\ell}f_{k}(x)e^{i\lambda^{\ell}a_{k}\cdot x}.$ (3.4)
Then, for any $M\in\mathbb{N}$ with $M>s$, there exists a positive constant $C$, depending only on $n$, $p$, $q$, $s$, $K$, and $M$ such that
${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant{}$ $C\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.5)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}$ (3.6)
@@
${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $C^{-1}\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\max_{1\leqslant k\leqslant K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.7)
$-C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-1}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}$ (3.8)
$-C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}$ (3.9)
@@
To obtain (3.6) with $q=1$, we may remove the directional assumption on $a_{1},\dots,a_{K}$. Indeed, applying (3.6) with $K=L=1$ to each $f_{k}(x)e^{i\lambda^{\ell}a_{k}\cdot x}$ gives
@@
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s}|b_{\ell}|{\left\|f_{k}\right\|}_{L^{p}}$ (3.11)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}.$ (3.12)
@@
$P_{m}\Delta_{j}g_{\lambda,\ell,k}$ $=e^{i\lambda^{\ell}a_{k}\cdot x}\mathscr{F}^{-1}\left[\widehat{\chi_{m,j}}(\xi+\lambda^{\ell}a_{k})\widehat{f_{k}}(\xi)\right]$ (3.16)
$=e^{i\lambda^{\ell}a_{k}\cdot x}\left(\chi_{m,j,\ell,k}*f_{k}\right)(x),$ (3.17)
@@
$\chi_{m,j}:=\mathscr{F}^{-1}[\phi_{m}\widehat{\varphi_{j}}],\qquad\chi_{m,j,\ell,k}(x):=e^{-i\lambda^{\ell}a_{k}\cdot x}\chi_{m,j}(x).$ (3.18)
For these kernels, there holds ${\left\|\chi_{m,j,\ell,k}\right\|}_{L^{\rho}}={\left\|\chi_{m,j}\right\|}_{L^{\rho}}=2^{n(1-\frac{1}{\rho})j}{\left\|\chi_{m,0}\right\|}_{L^{\rho}}$ holds for $1\leqslant\rho\leqslant\infty$. Moreover, we prepare some notations; for every $k\in\{1,\dots,K\}$ and $\ell\in\{1,\dots,L\}$, set
$j_{\ell,k}:=\log_{2}\left(\lambda^{\ell}|a_{k}|\right),$ (3.19)
$R_{\ell,k}:=\{(m,j)\in\{0,\dots,K\}\times\mathbb{Z}\ ;\ (m,j)\neq(k,j_{\ell,k}),(k,j_{\ell,k}\pm 1)\}.$ (3.20)
@@
There exists a constant $c_{*}=c_{*}(a_{1},\dots,a_{K})\in 2^{\mathbb{Z}}$ such that for any $\ell\in\{1,\dots,L\}$, $k\in\{1,\dots,K\}$, $(m,j)\in R_{\ell,k}$, and $\lambda\in 2^{\mathbb{N}}$, there holds
$\operatorname{supp}\left[\widehat{\chi_{m,j}}(\cdot+\lambda^{\ell}a_{k})\right]\subset\left\{\xi\in\mathbb{R}^{n}\ ;\ |\xi|\geqslant c_{*}\max\{2^{j},\lambda^{\ell}|a_{k}|\}\right\}.$ (3.21)
@@
Let $\xi\in\mathbb{R}^{n}$ satisfy $\widehat{\chi_{m,j}}(\xi+\lambda^{\ell}a_{k})\neq 0$. We first consider the case of $m=k$. Then, by $2^{j-1}\leqslant|\xi+\lambda^{\ell}a_{k}|\leqslant 2^{j+1}$ and $|\lambda^{\ell}a_{k}|=2^{j_{\ell,k}}$, it holds
@@
$\geqslant\begin{cases}|\xi+\lambda^{\ell}a_{k}|-|\lambda^{\ell}a_{k}|\geqslant 2^{j-1}-2^{j-2}&(j\geqslant j_{\ell,k}+2),\\ |\lambda^{\ell}a_{k}|-|\xi+\lambda^{\ell}a_{k}|\geqslant 2^{j_{\ell,k}}-2^{j_{\ell,k}-1}&(j\leqslant j_{\ell,k}-2),\end{cases}$ (3.23)
$\geqslant\frac{1}{4}\max\left\{2^{j},|\lambda^{\ell}a_{k}|\right\}.$ (3.24)
@@
$|\xi|^{2}={}$ $|\xi+\lambda^{\ell}a_{k}|^{2}+|\lambda^{\ell}a_{k}|^{2}-2|\xi+\lambda^{\ell}a_{k}||\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})$ (3.25)
$={}$ $\frac{1}{2}\left(|\xi+\lambda^{\ell}a_{k}|-|\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\right)^{2}$ (3.26)
$+\frac{1}{2}\left(|\lambda^{\ell}a_{k}|-|\xi+\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\right)^{2}$ (3.27)
$+\frac{1}{2}\sin^{2}\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\left(|\lambda^{\ell}a_{k}|^{2}+|\xi+\lambda^{\ell}a_{k}|^{2}\right)$ (3.28)
$\geqslant{}$ $\frac{1}{2}\sin^{2}(\vartheta_{*}/6)\left(|\lambda^{\ell}a_{k}|^{2}+|\xi+\lambda^{\ell}a_{k}|^{2}\right),$ (3.29)
which yields $|\xi|\geqslant c\sin(\vartheta_{*}/6)\max\left\{2^{j},|\lambda^{\ell}a_{k}|\right\}$. Thus, we complete the proof. ∎
@@
$F_{\lambda,L}={}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{k^{\prime}=0}^{K}\sum_{j\in\mathbb{Z}}P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}={}F_{\lambda,L}^{({\rm p})}+F_{\lambda,L}^{({\rm r})},$ (3.30)

### v2

@@ 1.3. Outline of the proof
Let us explain the idea of the proof of Theorem 1.2. The strategy is that we first establish a dominant part of the solution which lies in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with $(p,q)$ satisfying (N1) or (N2), and then solve the remaining perturbative problem by a fixed point argument in a function space where well-posedness is available. More precisely, we construct an almost frequency-localized building blocks $\{V_{j}\}_{j\geqslant 0}$ of stationary flows, each of which are essentially built on the $j$th frequency shell, and design the principal profile
$V=\sum_{j=0}^{\infty}V_{j}\in\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})\setminus\{0\}$ (1.11)
@@
$F:=-\Delta V+\mathbb{P}\operatorname{div}(V\otimes V)$ (1.12)
@@
${\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant C{\left\|F\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}\ll{\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}.$ (1.14)
Hence, the desired steady state is obtained in the form $U=V+W$. Note that $U\neq 0$ since
${\left\|U\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant{\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-C{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\geqslant\frac{1}{2}{\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}>0.$ (1.15)
Furthermore, to construct a non-stationary solution $u$ whose asymptotic profile is $U$, we decompose $u=U+w$. Then, we may follow the standard stability argument to construct a small dissipative perturbed flow $w\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some $2<\theta<\infty$. Then, the auxilliary dissipative space enables us to show $\|w(t)\|_{\dot{B}_{p,q}^{n/p-1}}\to 0$ as $t\to\infty$.
We now explain how to construct the leading profile $V$ satisfying (1.13). A natural first attempt, motivated by existing works such as [Coi-Pal-25], would be to choose each $V_{j}$ as a superposition of plane waves concentrated on a single highly lacunary frequency scale $N_{j}=A^{\gamma b^{j}}$ ($A\gg 1$), for instance, $V_{j}(x)=\sum_{k\in\mathcal{K}}f_{k,j}(x)\cos(N_{j}a_{k}\cdot x)k$, where $\mathcal{K}\subset\mathbb{R}^{n}$ is a finite set, $f_{k,j}$ are smooth amplitudes, and $a_{k}$ are constant vectors with $a_{k}\cdot k=0$. However, such a single-scale construction is too rigid to capture the delicate dependence on the Besov summability index $q$, which is essential for observing the subtle non-uniqueness phenomena addressed in (N1). Instead, we distribute the $j$th building block over many frequency scales. To this end, for $\lambda\gg 1$ (which is a parameter corresponding to $\varepsilon^{-1/\alpha}$ with some $\alpha$) and a fixed large number $\mu\in\mathbb{N}$, we introduce a tetration-like sequence [^4] This is defined by $\lambda_{0}=\lambda$ and $\lambda_{j}=\lambda^{2^{\lambda_{j-1}^{4\mu}}}$; see (4.9). $\{\lambda_{j}\}_{j\geqslant 0}$ and define the $j$th level frequency index set by
@@
and take $V_{j}$ in the schematic form [^5] In the actual construction of $V_{j}$, we apply a mollifier and introduce several additional modifications to enforce the divergence-free condition, so that its form becomes somewhat more involved.
@@
Here, the choice of $1/\sqrt{\ell}$ is inspired by [Yon-10, Tsu-19-ARMA]. With this multi-scale design, the oscillation calculus in Section 3 below yields
@@
The cross-shell interaction term $F_{2}$ is favorable, as it remains in high-frequency region. By contrast, $F_{1}$ exhibits a more delicate structure. Indeed, within the nonlinear interaction $\operatorname{div}(V_{j}\otimes V_{j})$, resonances among modes inside the frequency shell $\Lambda_{j}$ generate an inverse-cascade transfer toward low frequencies, which exhibit the worst regularity. By selecting the directional set $\mathcal{K}$ and the amplitudes $f_{k,j}$ appropriately through the Nash geometric lemma, one may arrange for this low-frequency contribution to cancel the linear part $\Delta V_{j-1}$. Consequently, the forcing term $F$ attains improved regularity and smaller magnitude relative to $V$.
@@ 2. Preliminaries
Throughout this paper, we denote by $C$ and $c$ the constants, which may differ in each line. In particular, C=C(,,)C=C(*,...,*) denotes the constant which depends only on the quantities appearing in parentheses. Let $\mathbb{N}$ be the set of all positive integers and put $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. For $\mathbb{M}\subset\mathbb{Z}$, we define $2^{\mathbb{M}}:=\{2^{m}\ ;\ m\in\mathbb{M}\}$. For two quantities $A_{+}$ and $A_{-}$, we use the notation $\sum_{\pm}A_{\pm}:=A_{+}+A_{-}$. The identity matrix is denoted by $\rm Id$}. For a metric space $(X,d)$, $x_{0}\in X$, and $r>0$, set $B_{X}(x_{0},r):=\{x\in X\ ;\ d(x,x_{0})\leqslant r\}$. Let $\mathscr{S}(\mathbb{R}^{n})$ be the set of all Schwartz functions on $\mathbb{R}^{n}$, and $\mathscr{S}^{\prime}(\mathbb{R}^{n})$ denotes the set of all tempered distributions on $\mathbb{R}^{n}$.
@@
$\Delta_{j}f:=\mathscr{F}^{-1}\left[\widehat{\varphi_{j}}(\xi)\widehat{f}(\xi)\right]$ (2.3)
for $f\in\mathscr{S}^{\prime}(\mathbb{R}^{n})$ and $j\in\mathbb{Z}$. For $1\leqslant p,q\leqslant\infty$ and $s\in\mathbb{R}$, the Besov space $\dot{B}_{p,q}^{s}(\mathbb{R}^{n})$ is defined as
@@
there exists a constant $C=C(n,p,q,r)>0$ such that
@@
The Chemin–Lerner spaces are well-known to be particularly convenient when exploiting the maximal regularity of the heat kernel. We take advantage of this framework through the following lemma, which will be used in our stability analysis for non-stationary flows around the steady state.
@@
Let $n\geqslant 2$. Let $p$, $q$, $\theta$ satisfy
$2\leqslant p<2n,\qquad 1\leqslant q\leqslant\infty,\qquad 2<\theta<\infty,\qquad\frac{n}{p}-1+\frac{1}{\theta}>0.$ (2.12)
Then, there exists a constant $C=C(n,p,q,\theta)>0$ such that
@@
for all $I=(t_{0},t_{1})\subset\mathbb{R}$, $v\in\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, and $w\in\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$.
@@ 3. Highly oscillating functions in Besov spaces
In our analysis, we frequently use the sharp Besov estimates for finite superpositions of functions obtained by modulating a smooth amplitude with a plane wave of the form (3.4) below. If each Fourier transform $\widehat{f_{k}}$ of the amplitudes were compactly supported in a low frequency region, then the Fourier support of each $g_{\lambda,k,\ell}$ is concentrated around $\xi\sim\lambda^{\ell}a_{k}$, and we have
${\left\|g_{\lambda,k,\ell}\right\|}_{\dot{B}_{p,q}^{s}}\sim|b_{\ell}|\lambda^{s\ell}{\left\|f_{k}\right\|}_{L^{p}}$ (3.1)
@@
$\|F_{\lambda,L}\|_{\dot{B}^{s}_{p,q}}\sim\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}{\left\|f_{k}\right\|}_{L^{p}}.$ (3.2)
@@
Let $K,L\in\mathbb{N}$. Let $a_{1},\dots,a_{K}\subset\mathbb{S}^{n-1}$ satisfy $a_{k}\neq a_{k^{\prime}}$ unless $k=k^{\prime}$. For $\lambda\in 2^{\mathbb{N}}$, $b_{1},\dots,b_{L}\in\mathbb{C}$, and $f_{1},\dots,f_{K}\in\mathscr{S}(\mathbb{R}^{n})$, set
$F_{\lambda,L}(x):=\sum_{k=1}^{K}\sum_{\ell=1}^{L}g_{\lambda,k,\ell}(x),\qquad g_{\lambda,k,\ell}(x):=b_{\ell}f_{k}(x)e^{i\lambda^{\ell}a_{k}\cdot x}.$ (3.4)
Then, for any $M\in\mathbb{N}$ with $M>s$, there exists a constant $C>0$, depending only on $n$, $p$, $q$, $s$, $K$, $a_{1},\dots,a_{K}$, and $M$ such that
${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant{}$ $C\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}{\left\|f_{k}\right\|}_{L^{p}}$ (3.5)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s+n(1-\frac{1}{p})-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}$ (3.6)
@@
${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $C^{-1}\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\max_{1\leqslant k\leqslant K}{\left\|f_{k}\right\|}_{L^{p}}$ (3.7)
$-C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s-1)\ell}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}$ (3.8)
$-C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s+n(1-\frac{1}{p})-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}$ (3.9)
@@
For (3.6) with $q=1$, we may remove the directional assumption on $a_{1},\dots,a_{K}$. Indeed, applying (3.6) with $K=L=1$ to each $g_{k,\ell}$ gives
@@
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{s\ell}|b_{\ell}|{\left\|f_{k}\right\|}_{L^{p}}$ (3.11)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s+n(1-\frac{1}{p})-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}.$ (3.12)
@@
$P_{m}\Delta_{j}g_{\lambda,k,\ell}$ $=e^{i\lambda^{\ell}a_{k}\cdot x}\mathscr{F}^{-1}\left[\widehat{\chi_{m,j}}(\xi+\lambda^{\ell}a_{k})\widehat{f_{k}}(\xi)\right]$ (3.16)
$=e^{i\lambda^{\ell}a_{k}\cdot x}\left(\chi_{m,j,k,\ell}*f_{k}\right)(x),$ (3.17)
@@
$\chi_{m,j}:=\mathscr{F}^{-1}[\phi_{m}\widehat{\varphi_{j}}],\qquad\chi_{m,j,k,\ell}(x):=e^{-i\lambda^{\ell}a_{k}\cdot x}\chi_{m,j}(x).$ (3.18)
For these kernels, there holds ${\left\|\chi_{m,j,k,\ell}\right\|}_{L^{\rho}}={\left\|\chi_{m,j}\right\|}_{L^{\rho}}=2^{n(1-\frac{1}{\rho})j}{\left\|\chi_{m,0}\right\|}_{L^{\rho}}$ holds for $1\leqslant\rho\leqslant\infty$. Moreover, we prepare some notations; for every $k\in\{1,\dots,K\}$ and $\ell\in\{1,\dots,L\}$, set
$j_{\ell}:=\ell\log_{2}\lambda,$ (3.19)
$R_{k,\ell}:=\{(m,j)\in\{0,\dots,K\}\times\mathbb{Z}\ ;\ (m,j)\neq(k,j_{\ell}),(k,j_{\ell}\pm 1)\}.$ (3.20)
@@
There exists a constant $c_{*}=c_{*}(K,a_{1},\dots,a_{K})\in 2^{\mathbb{Z}}$ such that for any $\ell\in\{1,\dots,L\}$, $k\in\{1,\dots,K\}$, $(m,j)\in R_{k,\ell}$, and $\lambda\in 2^{\mathbb{N}}$, there holds
$\operatorname{supp}\mathscr{F}\left[\chi_{m,j,k,\ell}\right]\subset\left\{\xi\in\mathbb{R}^{n}\ ;\ |\xi|\geqslant c_{*}\max\{2^{j},\lambda^{\ell}\}\right\}.$ (3.21)
@@
Let $\xi\in\mathbb{R}^{n}$ satisfy $\mathscr{F}\left[\chi_{m,j,k,\ell}\right](\xi)=\widehat{\chi_{m,j}}(\xi+\lambda^{\ell}a_{k})\neq 0$. We first consider the case of $m=k$. Then, by $2^{j-1}\leqslant|\xi+\lambda^{\ell}a_{k}|\leqslant 2^{j+1}$ and $\lambda^{\ell}=2^{j_{\ell}}$, it holds
@@
$\geqslant\begin{cases}|\xi+\lambda^{\ell}a_{k}|-\lambda^{\ell}\geqslant 2^{j-1}-2^{j-2}&(j\geqslant j_{\ell}+2),\\ \lambda^{\ell}-|\xi+\lambda^{\ell}a_{k}|\geqslant 2^{j_{\ell}}-2^{j_{\ell}-1}&(j\leqslant j_{\ell}-2),\end{cases}$ (3.23)
$\geqslant c\max\left\{2^{j},\lambda^{\ell}\right\}.$ (3.24)
@@
$|\xi|^{2}={}$ $|\xi+\lambda^{\ell}a_{k}|^{2}+\lambda^{2\ell}-2\lambda^{\ell}|\xi+\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})$ (3.25)
$={}$ $\frac{1}{2}\left(|\xi+\lambda^{\ell}a_{k}|-\lambda^{\ell}\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\right)^{2}$ (3.26)
$+\frac{1}{2}\left(\lambda^{\ell}-|\xi+\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\right)^{2}$ (3.27)
$+\frac{1}{2}\sin^{2}\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\left(\lambda^{2\ell}+|\xi+\lambda^{\ell}a_{k}|^{2}\right)$ (3.28)
$\geqslant{}$ $\frac{1}{2}\sin^{2}(\vartheta_{*}/6)\left(\lambda^{2\ell}+|\xi+\lambda^{\ell}a_{k}|^{2}\right),$ (3.29)
which yields $|\xi|\geqslant c\max\left\{2^{j},\lambda^{\ell}\right\}$. Thus, we complete the proof. ∎
@@
$F_{\lambda,L}={}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{m=0}^{K}\sum_{j\in\mathbb{Z}}P_{m}\Delta_{j}g_{\lambda,k,\ell}={}F_{\lambda,L}^{({\rm p})}+F_{\lambda,L}^{({\rm r})},$ (3.30)